I also need help solving this literal equations. I need to solve for x.
cx + 4c = 5x + 7c

Answer by Edwin McCravy(20081)

                                                                                             
                                              We only want x terms on the left, 
                         cx + 4c = 5x + 7c    so get the x term off the right
                   -5x            -5x         by adding -5x to both sides.
                   ------------------------
                   -5x + cx + 4c =      7c    4c does not contain x, so 
                            - 4c      - 4c    get if off the left side by
                   ------------------------   adding -4c to both sides.
                   -5x + cx      =      3c 
                                              Factor out x on the left
                  x(-5 + c)      =      3c
                                              Get rid of the multiplied -5+c
                                              by dividing both sides by it
                  x(-5 + c)             3c
                 覧覧覧覧覧      =   覧覧覧覧覧
                   -5 + c             -5 + c             
                            
                                              Cancel the -5+c on the left:

                                        3c
                         x       =   覧覧覧覧覧
                                      -5 + c 

That's the answer, but maybe your teacher would think it would look better
if you wrote the denominator like this:

                                        3c
                         x       =   覧覧覧覧覧
                                       c - 5 

Edwin